OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a base(1) = 16 and a base(15) = -40 ?

OpenStudy (anonymous):

a(n) = a(1) + d(n - 1) -4 = 16 + d(16 - 1) -40 = 16 + 15d -56 = 15d d = -3.73? This is where I'm stuck. I understand that the next formula is: a(n) = a(1) + d(n - 1) But is -3.73 correct?

OpenStudy (anonymous):

-40= 16 + d(15 - 1) so d=-4

OpenStudy (anonymous):

a[41) =-144